Assume that Mr. Binda is a speculator who buys a 90-day British Pound call option with a strike price of $27.02. Assume one option contract specifies 10,000 units and the current spot price as of that date is $26.87. Mr. Binda pays a premium of $0.05 per unit for the call option and no other charge (such as brokerage fee). Just on the expiration date, the spot rate of a pound reaches $27.18.

Required:-

a) Determine the profit or loss if the option is exercised,

b) Determine the value of the call option lithe option is exercised.



1
Expert's answer
2021-03-23T08:24:47-0400

initial price

"a) 10000\times27.02=\$272,000"

price after call option "=10000\times26.87-0.05\times10000=\$268,200"

profit"=272,000-268,200=\$3800"

b) "10000\times27.18=\$271,800"